# Pay Attention: a Call to Regulate the Attention Market and Prevent Algorithmic Emotional Governance

### v1

@@ 4 Attention, attention, all thinkers
We firstly intended this section for all the scientists reading this paper, concurring with the article of David R. Smith: “Attention, attention: your most valuable scientific assets are under attack” [60]. In this article, Smith calls for attention to what media platforms are doing to research and the academic domain. Indeed, even the most informed scientists and engineers are not immune to these problems [37] such that digital contraptions (as Smith calls them) are contributing to academic attention deficit disorder [60]. In fact, concentration but also boredom, mind-wandering and daydreaming times are vital to creative thinking. Many of us experienced the sudden burst of an idea in the middle of a relaxing moment. Attention-capturing systems steal these moments from all of us and hamper the creativity process of wondering minds [77].
Of course these remarks can be generalized to many other activities requiring concentration, creativity and imagination, and one could wonder what digital contraptions are doing to politics, healthcare or education, for instance. To mention just one example, countless information media now report the cases of Youtubers experiencing a burnout [46], or musicians complaining that they spend more time making Tiktok videos to promote their music than actually creating music [57, 73]. This reveals that, to hook and keep the attention of content consumers, platforms also exercise some sort of visibility tyranny over content creators.
In other words: attention, attention, thinkers, we need to redesign the systems for our own needs, rather than the other way around, especially in creative jobs since the true currency of these jobs are ideas [60].
@@ 5 Known Post Hoc Counter-measures
Among the various issues raised in the previous sections, the questions of false information, radicalization, hateful speech and bullying are among the most concerning, and therefore have been extensively addressed by the research community [58]. In [22] authors identified three different points where recommendation systems can be adapted to tackle these issues: 1 pre recommendation, 2 within the recommendation model, and 3 post recommendation. Most of the current counter-measures to deal with false information lie in this third category. Below we touch upon some of them.
@@
Firstly, to dyke the spread of false information as well as inappropriate content such as bullying, hateful or violent speech, social media and content hosting platforms have obligations that vary depending on the legislation and its jurisdiction [24]. Measures range from content deletion and suspension of users spreading inappropriate content, to re-ranking of recommended items before presenting them to the user [22], flagging to indicate potentially deceptive content, etc. Yet, despite these various approaches, progress is still necessary. For instance, subtle violent content may be hard to detect as soon as it does not contain explicit hateful or violent terms, or when it uses sarcasm [44]. Conversely, content may be erroneously assessed as abusive or illicit although it is in fact using irony to convey perfectly acceptable ideas. An in-depth analysis of implicit and subtlety in linguistic content remains an open question [44].
@@
In addition, any action must carefully consider the dangers of transferring regulation and enforcement to private companies.  [62] argue that over-filtering content is just as dangerous as letting bad content spread. Indeed, deletion and filtering may deviate from initial purpose to over-censorship of content if it becomes safer for the platforms to do so than take a risk of being sued. Furthermore, assessing the trustworthiness of information raises multiple ethical and political concerns: Who decides what is true or false? According to which criteria? Under whose control?
@@
Secondly, to mitigate the effect of false information, multiple post hoc measures rely on the impact of additional corrective content. For [70], pointing to a coherent alternative explanation, with references to expert and highly credible factual sources, remains a solid starting point. The authors describe the strategy of “observational correction” leveraging the fact that users who witness the correction of a misinformation item, but have not directly engaged with that item, are less affected by cognitive dissonance and are therefore more amenable to correction. This is consistent with the findings of [8] who suggest that exposing users to related stories that correct a post that contained misinformation will significantly reduce misperceptions. The impact of the correction can be further reinforced by explicitly pointing to the demographic similarity between the user and the authors of opposing content [25], which taps into the homophily effect [^10] Homophily: the tendency to associate with similar others.. In other words, we are more likely to accept the correction when it comes from someone who is socially close to us, e.g. having the same professional activity or background. [70] also suggest to multiply correction actions for each targeted content to reinforce the effect.
@@ 6 Reclaiming our attention
The methods presented in the previous section all have one thing in common: they deal with the problems in a post hoc manner, that is, after these problems have occurred, with all the limitations that come with this “coming after”. To go further however, we need to figure out measures, may they be legal, political or technical, capable of preventing the attention from being looted in the first place. More importantly, we need to consider this reflection not only from the perspective of regulating the attention consumers (the platforms and multiple intermediaries), but also from the perspective of the producers (the end-users) who want to reclaim their attention, especially in times when our attention is needed on a number of urgent matters. This involves actively preventing recommendation systems from finding ways to exploit our inner limitations and manipulate us through sometime ancient and deeply embedded structures of our brain (e.g. our striatum) [9].
@@
6.1 The carrot and the stick
Taking the example of false information, [70] insist on the fact that a posteriori corrections are not sufficient and must happen as early as possible, that is, before misperceptions are entrenched. Besides, avoiding the algorithmic amplification effect of such information by recommendation systems requires to mitigate the popularity effect before its happens [22]. But if online social media are required by law to combat false information, they have conflicting incentive to do it, not to say no incentive at all. Indeed, as we described in section 3, false information largely relies on negative emotions to capture users’ attention. As such, they are very effective in fostering user engagement which is what online social media strive to obtain. Consequently, from an economic point of view, it is counter-productive for online social media to prevent the spread of false information. More generally, it is counter-productive for platforms to mitigate the popularity effect, mitigate the impact of negative emotions, or reduce filter bubbles and the subsequent polarization of opinions.
The authors of [42] suggest to rethink existing trade regulation laws such as antitrust and fair competition laws under the new realm of attention markets. They propose to enforce taxes on attention consumption to “disincentivize attention intermediaries from vacuuming up as much attention as possible”, for instance by restraining the amount of advertisements that can be shown to a user, or reducing the deductibility of advertising expenditures from the companies’ revenues to alleviate their taxes. They also propose to regulate the attention costs that can be charged, with the idea that if attention becomes less lucrative then financial resources will be redirected towards more lucrative markets, thus reducing the amount of attention being captured and traded.
@@ 6.1 The carrot and the stick
At a Web governance level, we must leverage legal and economic means to drive platforms’ practices towards desirable behaviours, while penalizing undesired behaviours.
6.2 Usage regulation
Some of the measures meant to regulate the attention market lie in the way the services are consumed. Some legal measures could be taken to enforce a regulation of the daily use of Web platforms. As has already been done in some countries, laws could be voted to limit the daily time spent by users on certain services, especially among the youngest [32, 11]. Another simple measure applies to video streaming platforms, that consists in imposing few-second pauses between videos. This apparently naive technique may actually shatter the infinite feed trap by giving users the short amount of time they need to realize that they have been in an attention tunnel for a while, and that they want to “reclaim” their attention. This can be generalized by formulating the following principle:
@@ 6.2 Usage regulation
Policy makers could also tackle the problem of attention fragmentation entailed by the multiple, often invasive, notifications that smartphone applications raise. Whenever a notification occurs, users are tempted to interrupt their current activity, check the reason of the notification, possibly react to it, before eventually returning to their activity. It has been shown that switching our attention between tasks or contexts has a cost: it is time-consuming and creates a more error prone context [31, 40, 52]. Furthermore it has even been shown that the mere presence of such devices, although turned off, impairs our cognitive capacity [71]. In a way similar to the European General Data Protection Regulation (GDPR) [^11] General Data Protection Regulation (GDPR) https://gdpr-info.eu/ which imposes the consent of users for the use of cookies, law could impose that smartphone applications obtain users’ explicit and informed consent for the notifications that they raise, and deactivate them by default (“opt-in only”). Hence the following principle:
@@
This could be complemented by more punitive measures, as proposed in [41], for instance by demonetizing and forbidding collaboration with platforms that do not follow the rules.
6.3 Content recommendation monitoring
The echo-chamber effect of recommendation algorithms is at the root of multiple examples of polarization and radicalization. It could be mitigated by imposing a certain share of non-recommended content, content that is outside of the user’s interests, or content that originates from users they are not acquainted with. In this respect, some approaches lie in the second category proposed by  [22], i.e. modifications “within the recommendation” system. For instance, the same authors suggest using clustering approaches to assemble the contacts of the user according to different levels of similarity with the user, and leverage these groups to increase the diversity of recommendations while maintaining a certain coherence and similarity. [20] propose a method to come up with relevant recommendations while reducing the likelihood of enticing the user towards radicalization pathways. Also, to counter the misuse and abuse of anger, indignation or fear, which are often associated with false information, platforms could be required to carry out sentiment analysis on every content in order to keep the amount of recommendations associated to negative emotions below a given threshold.
@@ 6.3 Content recommendation monitoring
Moreover, there exists an asymmetry of visibility between the viral spreading of an information that was proven to be false or misleading, and the denial or rectification of that information. The denial of a false information usually puts forward a pondered, nuanced position that appeals to reasoning and facts (logos) over emotion (pathos). Hence, it does not trigger an emotional response compared to the one generated by the false information in the first place, and it is therefore silently downgraded by recommendation algorithms. This is commonly summarized by the so-called Bradolini’s law which states that “the amount of energy needed to refute false information is an order of magnitude bigger than that needed to produce it.” As a result, users who propagate false information often never get to know about their mistake.  [67] insist on the fact that it is critical to jointly address content-checking and digital virality. Thus, to counterbalance this visibility asymmetry, social media could be required to impose on the denial/rectification of a false information a visibility equivalent to that of the initial information, for instance, by ensuring that the population who was exposed to the false information be exposed to the denial too. A warning could also be presented to users who propagated this false information in order to increase their awareness. Of course, this type of measure could be coupled with other post recommendation measures such as strategies involving the observational correction or demographic similarity presented in section 5. More generally, one has to figure out how we can use recommendation systems to recommend counter measures, i.e. we could train a recommendation system to learn the most relevant content and the most impactful entries in the acquaintance network to inject a counter measure.
@@
6.4 Affordances and interaction design
@@
6.5 Societal impact and educational mission
We, as a society, could decide that large online social media, because of their influence on the society, public opinion, public health and economy, can no longer be considered as sheer private companies regulated by markets law only. Instead, they could be seen as digital commons and be assigned a specific status that would endow them with a societal mission including an educational purpose, for instance. As an example, they could instruct users in detecting false or misleading information, they could promote content meant to increase awareness w.r.t. attention mechanisms and cognitive biases, foster critical thinking and “distill” the scientific method, etc. On the same page, authors of [1] insist on the need for civic education, and [43] recommend integrating democratic values into the algorithms that impact our lives, especially the ones participating in an algorithmic governance (e.g. platform used for debates, for information, for legal actions, educational orientation) which, in our case, means going beyond the optimization of user engagement and attention catching, and including ethical criteria.
@@ 6.5 Societal impact and educational mission
6.6 Feedback and transparency enforcement
Since one of the pitfalls we identified is the fact that users are being caught by the recommendation loops, approaches such as quantified self and lifelog could be specialized to the case of recommendation-based platforms, in order to foster awareness and introspection. Self-tracking tools could provide users with usage metrics and feedback with respect to the total time spent on the platform, the total exposure to negative news, etc. This could be a way to counter the fact that online sharing of fake news increases with social media fatigue [64]. Indicators could inform users about the diversity of the recommendations they are shown, and make them aware of low-diversity risks. For instance the fact that “90% of the content one sees come from 10% of one’s contacts or are on the same topic” may indicate that one is experiencing a filter bubble. We summarize this in the principle below where user’s reflexivity is the ability of users to be self-aware of their usage and engagement with the system.
@@ 6.6 Feedback and transparency enforcement
Among other measures, the European Digital Services Act [66], that took effect in August 2023, requires that platforms set up mechanisms to explain the reasons that led to recommending a certain content, and to offer users an alternative recommendation not based on profiling. Such measures are especially crucial when coupling AI and the Web since we need to set transparency and explanation as a prerequisite to any approach, to ensure the awareness and informed consent of, potentially, billions of users [6].
@@
6.7 Build on existing practices
@@
Methods and tools used to regulate similar situations in relevant domains should be surveyed, benchmarked and systematically considered as input to a Web governance.
To give another example coming from a completely different angle, we know that parenting practices in terms of TV viewing have an impact on the behaviour of young watchers [3]. Again, approaches and good practices in this domain, and more generally in educating and parenting in the digital media age [16], must be considered in the case of “Web viewing” in general and when addressing the problem of attention capturing in particular.
More generally speaking, we need to put in place a governance bodies, starting with the Web and AI, that are prepared to tackle new problematic practices and regulate them, as is done in other areas of activity. And we also need to keep a constant watch on these other areas, if only to draw inspiration from the initiatives and feedback they have on similar issues. Taking the example of the video game industry, there is evidence of a relationship between “loot box”141414“loot boxes” are video game items with randomized contents that can be paid for with real-world money. spending and gambling addiction [76], and that a loot box is psychologically akin to gambling [19] and can result in addictive behaviors and endangered players. The way to study and address that unwanted exploitation of our behaviors is inspirational for other problematic practices on the Web such as those we surveyed.

### v2

@@ 4 Attention, Attention, all Thinkers
We firstly intended this section for all the scientists reading this paper, concurring with the article of David R. Smith: “Attention, attention: your most valuable scientific assets are under attack” (Smith 2018). In this article, Smith calls for attention to what media platforms are doing to research and the academic domain. Indeed, even the most informed scientists and engineers are not immune to these problems (Lewis 2017) such that digital contraptions (as Smith calls them) are contributing to academic attention deficit disorder (Smith 2018). In fact, concentration but also boredom, mind-wandering and daydreaming times are vital to creative thinking. Many of us experienced the sudden burst of an idea in the middle of a relaxing moment. Attention-capturing systems steal these moments from all of us and hamper the creativity process of wondering minds (Zomorodi 2017).
Of course these remarks can be generalized to many other activities requiring concentration, creativity and imagination, and one could wonder what digital contraptions are doing to politics, healthcare or education, for instance. To mention just one example, countless information media now report the cases of Youtubers experiencing a burnout (Parkin 2018), or musicians complaining that they spend more time making Tiktok videos to promote their music than actually creating music (Shah 2022; Whateley 2023). This reveals that, to hook and keep the attention of content consumers, platforms also exercise some sort of visibility tyranny over content creators.
In other words: attention, attention, thinkers, we need to redesign the systems for our own needs, rather than the other way around, especially in creative jobs since the true currency of these jobs are ideas (Smith 2018).
@@ 5 Known Post Hoc Counter-Measures
Among the various issues raised in the previous sections, the questions of false information, radicalization, hateful speech and bullying are among the most concerning, and therefore have been extensively addressed by the research community (Sharma et al. 2019). In (Fernández, Bellogín, and Cantador 2021) authors identified three different points where recommendation systems can be adapted to tackle these issues: 1 pre recommendation, 2 within the recommendation model, and 3 post recommendation. Most of the current counter-measures to deal with false information lie in this third category. Below we touch upon some of them.
@@
Firstly, to dyke the spread of false information as well as inappropriate content such as bullying, hateful or violent speech, social media and content hosting platforms have obligations that vary depending on the legislation and its jurisdiction (Funke and Flamini 2023). Measures range from content deletion and suspension of users spreading inappropriate content, to re-ranking of recommended items before presenting them to the user (Fernández, Bellogín, and Cantador 2021), flagging to indicate potentially deceptive content, etc. Yet, despite these various approaches, progress is still necessary. For instance, subtle violent content may be hard to detect as soon as it does not contain explicit hateful or violent terms, or when it uses sarcasm (Ocampo et al. 2023). Conversely, content may be erroneously assessed as abusive or illicit although it is in fact using irony to convey perfectly acceptable ideas. An in-depth analysis of implicit and subtlety in linguistic content remains an open question (Ocampo et al. 2023).
@@
In addition, any action must carefully consider the dangers of transferring regulation and enforcement to private companies.  (Stark et al. 2020) argue that over-filtering content is just as dangerous as letting bad content spread. Indeed, deletion and filtering may deviate from initial purpose to over-censorship of content if it becomes safer for the platforms to do so than take a risk of being sued. Furthermore, assessing the trustworthiness of information raises multiple ethical and political concerns: Who decides what is true or false? According to which criteria? Under whose control?
@@
Secondly, to mitigate the effect of false information, multiple post hoc measures rely on the impact of additional corrective content. For (Vraga and Bode 2020), pointing to a coherent alternative explanation, with references to expert and highly credible factual sources, remains a solid starting point. The authors describe the strategy of “observational correction” leveraging the fact that users who witness the correction of a misinformation item, but have not directly engaged with that item, are less affected by cognitive dissonance and are therefore more amenable to correction. This is consistent with the findings of (Bode and Vraga 2015) who suggest that exposing users to related stories that correct a post that contained misinformation will significantly reduce misperceptions. The impact of the correction can be further reinforced by explicitly pointing to the demographic similarity between the user and the authors of opposing content (Garrett, Nisbet, and Lynch 2013), which taps into the homophily effect [^8] Homophily: the tendency to associate with similar others.. In other words, we are more likely to accept the correction when it comes from someone who is socially close to us, e.g. having the same professional activity or background. (Vraga and Bode 2020) also suggest to multiply correction actions for each targeted content to reinforce the effect.
@@ 6 Reclaiming Our Attention
The methods presented in the previous section all have one thing in common: they deal with the problems in a post hoc manner, that is, after these problems have occurred, with all the limitations that come with this “coming after”. To go further however, we need to figure out measures, may they be legal, economical or technical, capable of preventing the attention from being looted in the first place. More importantly, we need to consider this reflection not only from the perspective of regulating the attention consumers (the platforms and multiple intermediaries), but also from the perspective of the producers (the end-users) who want to reclaim their attention, especially in times when our attention is needed on a number of urgent matters. This involves actively preventing recommendation systems from finding ways to exploit our inner limitations and manipulate us through sometime ancient and deeply embedded structures of our brain (e.g. our striatum) (Bohler 2019).
@@
The Carrot and the Stick
Taking the example of false information, (Vraga and Bode 2020) insist on the fact that a posteriori corrections are not sufficient and must happen as early as possible, that is, before misperceptions are entrenched. Besides, avoiding the algorithmic amplification effect of such information by recommendation systems requires to mitigate the popularity effect before its happens (Fernández, Bellogín, and Cantador 2021). But if online social media are required by law to combat false information, they have conflicting incentive to do it, not to say no incentive at all. Indeed, as we described in section 3, false information largely relies on negative emotions to capture users’ attention. As such, they are very effective in fostering user engagement which is what online social media strive to obtain. Consequently, from an economic point of view, it is counter-productive for them to prevent the spread of false information. More generally, it is counter-productive for platforms to mitigate the popularity effect, mitigate the impact of negative emotions, or reduce filter bubbles and the subsequent polarization of opinions.
The authors of (Newman 2019) suggest to rethink existing trade regulation laws such as antitrust and fair competition laws under the new realm of attention markets. They propose to enforce taxes on attention consumption to “disincentivize attention intermediaries from vacuuming up as much attention as possible”, for instance by restraining the amount of advertisements that can be shown to a user, or reducing the deductibility of advertising expenditures from the companies’ revenues to alleviate their taxes. They also propose to regulate the attention costs that can be charged, with the idea that if attention becomes less lucrative then financial resources will be redirected towards more lucrative markets, thus reducing the amount of attention being captured and traded. Such regulation could take the form of a tax on the daily attention time being caught beyond a certain threshold. Another tax could be enforced on the individual and collective harms caused by attention capturing techniques, which first requires to explicit those harms like it has been done for dark patterns (Santos, Morozovaite, and De Conca 2024) and more precisely to qualify (Beckers, Chockler, and Halpern 2022) and quantity (Beckers, Chockler, and Halpern 2023) the harms done.
@@
Usage Regulation
Some of the measures meant to regulate the attention market lie in the way the services are consumed. As has already been done in some countries, laws could be voted to limit the daily time spent by users on certain services, especially among the youngest (Kantrowitz 2021; Carroll and correspondent 2023). Another simple measure applies to video streaming platforms, that consists in imposing few-second pauses between videos. This apparently naive technique may actually shatter the infinite feed trap by giving users the short amount of time they need to realize that they have been in an attention tunnel for a while, and that they want to “reclaim” their attention. This can be generalized by formulating the following principle:
@@
Policy makers could also tackle the problem of attention fragmentation entailed by the multiple, often invasive, notifications that smartphone applications raise. Whenever a notification occurs, users are tempted to interrupt their current activity, check the reason of the notification, possibly react to it, before eventually returning to their activity. It has been shown that switching our attention between tasks or contexts has a cost: it is time-consuming and creates a more error prone context (Jersild 1927; Monsell 2003; Rogers and Monsell 1995). Furthermore it has even been shown that the mere presence of such devices, although turned off, impairs our cognitive capacity (Ward et al. 2017). In a mindset similar to the European General Data Protection Regulation (GDPR) [^9] General Data Protection Regulation (GDPR)
https://gdpr-info.eu/
, which imposes the consent of users for the use of cookies, the EU Parliament has recently adopted a resolution on addictive design to put forward “a digital ‘right not to be disturbed’ to empower consumers by turning all attention-seeking features off by design” (The European Parliament 2023). This could lead to laws imposing that smartphone applications obtain users’ explicit and informed consent for the notifications that they raise, and deactivate them by default (“opt-in only”). Hence the following principle:
@@
This could be complemented by more punitive measures, as proposed in (Mudde and Kaltwasser 2017), for instance by demonetizing and forbidding collaboration with platforms that do not follow the rules.
Content Recommendation Monitoring
The echo-chamber effect of recommendation algorithms is at the root of multiple examples of polarization and radicalization. It could be mitigated by imposing a certain share of non-recommended content, content that is outside of the user’s interests, or content that originates from users they are not acquainted with. In this respect, some approaches lie in the second category proposed by  (Fernández, Bellogín, and Cantador 2021), i.e. modifications “within the recommendation” system. For instance, the same authors suggest using clustering approaches to assemble the contacts of the user according to different levels of similarity with the user, and leverage these groups to increase the diversity of recommendations while maintaining a certain coherence and similarity. (Fabbri et al. 2022) propose a method to come up with relevant recommendations while reducing the likelihood of enticing the user towards radicalization pathways. Also, to counter the misuse and abuse of anger, indignation or fear, which are often associated with false information, platforms could be required to carry out sentiment analysis on every content in order to keep the amount of recommendations associated to negative emotions below a given threshold.
@@
Moreover, there exists an asymmetry of visibility between the viral spreading of an information that was proven to be false or misleading, and the denial or rectification of that information. The denial of a false information usually puts forward a pondered, nuanced position that appeals to reasoning and facts (logos) over emotion (pathos). Hence, it does not trigger an emotional response compared to the one generated by the false information in the first place, and it is therefore silently downgraded by recommendation algorithms. This is commonly summarized by the so-called Bradolini’s law which states that “the amount of energy needed to refute false information is an order of magnitude bigger than that needed to produce it.” As a result, users who propagate false information often never get to know about their mistake.  (Venturini 2019) insist on the fact that it is critical to jointly address content-checking and digital virality. Thus, to counterbalance this visibility asymmetry, social media could be required to impose on the denial/rectification of a false information a visibility equivalent to that of the initial information, for instance, by ensuring that the population who was exposed to the false information be exposed to the denial too. A warning could also be presented to users who propagated this false information in order to increase their awareness. Of course, this type of measure could be coupled with other post recommendation measures such as strategies involving the observational correction or demographic similarity presented in section 5. More generally, one has to figure out how we can use recommendation systems to recommend counter measures, i.e. we could train a recommendation system to learn the most relevant content and the most impactful entries in the acquaintance network to inject a counter measure.
@@
Affordances and Interaction Design
@@
Societal Impact and Educational Mission
We, as a society, could decide that large online social media, because of their influence on the society, public opinion, public health and economy, can no longer be considered as sheer private companies regulated by markets law only. Instead, they could be seen as digital commons and be assigned a specific status that would endow them with a societal mission including, for instance, an educational purpose. As an example, they could instruct users in detecting false or misleading information, they could promote content meant to increase awareness w.r.t. attention mechanisms and cognitive biases, foster assertiveness, critical thinking and “distill” concepts of the scientific method, etc. On the same page, authors of (Abreu and França 2021) insist on the need for civic education, and (O’neil 2017) recommend integrating democratic values into the algorithms that impact our lives, especially the ones participating in an algorithmic governance (e.g. platform used for debates, for information, for legal actions, education). States and civil societies could favor the entrance of new platforms that, instead of optimizing user engagement and attention catching, would favor democratic debate in the digital space by supporting free constructive debate and the collective development of solutions to major societal issues. An illustration of this approach is that of Polis, an open-source social network meant to support large-scale public debate (Small et al. 2021). Its recommendation algorithm is trained to promote consensus. During experimentations, particularly in Taiwan, Polis has “shown itself capable of building shared understanding, disincentivizing counterproductive behavior (trolling), and cultivating points of consensus”.
@@
Feedback and Transparency Enforcement
One of the pitfalls we identified is the fact that users are not aware of being caught by recommendation loops. Furthermore, it is established that online sharing of fake news increases with social media fatigue (Talwar et al. 2019). Therefore, approaches such as quantified self and lifelog could be specialized to the case of recommendation-based platforms, in order to foster awareness and introspection. Some online services already provide their users with usage metrics with respect to the total time spent on the platform, or reminders to make a break at regular intervals. This can be seen as a shy improvement towards compliance with the European Digital Services Act (Union 2022), that took effect in August 2023, however such options remain deactivated by default, and usually accessible from impractical sub-menus. Self-tracking tools could provide users with a much larger set of metrics, indicators and feedback such as the total exposure to negative news, the diversity of the recommendations they are shown, and make them aware of low-diversity risks. For instance the fact that “90% of the content one sees come from 10% of one’s contacts or are on the same topic” may indicate that one is experiencing a filter bubble. In other words, platforms must set up the conditions for individual and collective reflexivity, that is, the ability of users and communities to be self-aware of their usage and engagement with the system. Hence the following principle:
@@
This reflexivity requires not only usage metrics but also transparency about the platform’s purposes. Among other measures, the European Digital Services Act requires that platforms set up mechanisms to explain the reasons that led to recommending a certain content, and to offer users an alternative recommendation not based on profiling. Such measures are especially crucial when coupling AI and the Web since we need to set transparency and explanation as a prerequisite to any approach, to ensure the awareness and informed consent of, potentially, billions of users (Berendt et al. 2021).
@@
Let us add that, for the regulations proposed above to be enforced effectively, one cannot rely solely on the goodwill of online platforms. But given the scale of these platforms, public actors have neither the information nor the technical possibilities to evaluate the application and effectiveness of the regulations (Zolynski, Le Roy, and Levin 2019). Hence the following principle:
[…]
To the very least, this requires open and free access to platforms data by researchers, as opposed to the recent decision of X to deny them free API access.
Build on Existing Practices
@@
To give another example coming from a completely different angle, we know that parenting practices in terms of TV viewing have an impact on the behaviour of young watchers (Barradas et al. 2007). Again, approaches and good practices in this domain, and more generally in educating and parenting in the digital media age (Coyne et al. 2017), must be considered in the case of “Web viewing” in general and when addressing the problem of attention capturing in particular.
More generally speaking, we need to put in place a governance bodies, starting with the Web and AI, that are prepared to tackle new problematic practices and regulate them, as is done in other areas of activity. And we also need to keep a constant watch on these other areas, if only to draw inspiration from the initiatives and feedback they have on similar issues. Taking the example of the video game industry, there is evidence of a relationship between “loot box”121212“loot boxes” are video game items with randomized contents that can be paid for with real-world money. spending and gambling addiction (Zendle and Cairns 2018), and that a loot box is psychologically akin to gambling (Drummond and Sauer 2018) and can result in addictive behaviors and endangered players. The way to study and address that unwanted exploitation of our behaviors is inspirational for other problematic practices on the Web such as those we surveyed.